# Muonium HFS Uncertainty Revisited

### v1

@@ Abstract
Uncertainty of the quantum electrodynamics theoretical prediction for the hyperfine splitting in the ground state of muonium is considered. It is compared with the respective discussion in the two most recent CODATA adjustments of the fundamental physical constants.
I Introduction
High precision measurements of HFS in muonium for a long time were considered as a test of the high precision QED and a source for precise values of the fine structure constant $\alpha$ and the muon-electron mass ratio $m_{\mu}/m_{e}$ [2, 3, 1]. While the role of muonium HFS in determining the fine structure constant was made obsolete by the highly precise $\alpha$ obtained from the measurements of the electron anomalous magnetic moment $a_{e}$ [4] and the recoil frequency of the $^{133}Cs$ [5] and $^{87}Rb$ [6] atoms, it remains the best source for the precise value of the muon-electron mass ratio. The hyperfine splitting in muonium and muon-electron mass ratio were measured some time ago in two LAMPF experiments [7, 8]
$\Delta\nu^{ex}_{\scriptscriptstyle HFS}(Mu)=4~463~302~776~(51)~\mbox{Hz},\quad\delta=1.1\times 10^{-8},$ (1)
@@ I Introduction
$\Delta\nu_{\scriptscriptstyle weak}$ is the $Z$-boson exchange contribution [9, 10, 11, 12], $\Delta\nu_{\scriptscriptstyle hadr}$ is the hadron vacuum polarization contribution [13, 14, 15] and $\Delta\nu_{th}$ is the estimate of the yet uncalculated terms, see, e.g., [2, 3]. Numerically [16] [^1] This is an updated value of HFS, which takes into account some minor changes in theoretical contributions and values of the fundamental constants, which happened after [16] was published. Only the last digit changed by one, what does not change any conclusions of [16].
$\Delta\nu^{th}_{\scriptscriptstyle HFS}(Mu)=4~463~302~873~(511)~(70)~(2)~\mbox{Hz},$ (5)
@@
$\Delta\nu^{th}_{\scriptscriptstyle HFS}(Mu)=4~463~302~873~(515)~\mbox{Hz},\quad\delta=1.2\times 10^{-7}.$ (6)
@@
$\Delta\nu_{\scriptscriptstyle Mu}(th)+\delta_{\scriptscriptstyle th}(Mu)=4~463~302~776~(51)~{\rm Hz}~~~~[1.1\times 10^{-8}]$ (7)
This “recommended value for the muonium hyperfine splitting” is never defined in the text but its value and uncertainty precisely coincide with the value and uncertainty of the experimental HFS in Eq. (1). This is not a surprise, if we assume that the experimental HFS was treated in [21, 22] as a fundamental constant and was included in the least square adjustment. As we already mentioned, the last precise experiment on muonium HFS was done in 1999 [7, 8], and up to this moment there are no comparable in accuracy sources for the experimental value of the muonium HFS. This means that no adjustment can significantly change the values and uncertainties of muonium HFS. However, the very idea of including the experimental HFS in the least square adjustment seems to be dubious.
But this is not how “the recommended value for the muonium hyperfine splitting” is treated in [21, 22]. By appearance, “the recommended value for the muonium hyperfine splitting” [21, 22] in Eq. (7) looks like a theoretical QED prediction, it is called $\Delta E_{\scriptscriptstyle Mu}(th)+\delta_{\scriptscriptstyle th}(Mu)$. Moreover, “the recommended value for the muonium hyperfine splitting” is explicitly interpreted in [21, 22] as a theoretical QED prediction, it is compared with the theoretical QED prediction in [16]. The authors of [21, 22] write: “Eides (2019) gave an alternative prediction for the uncertainty of the recommended muonium hyperfine splitting.” This statement demonstrates the “the recommended muonium hyperfine splitting” is considered in [21, 22] as a theoretical QED prediction for muonium HFS since this is what was discussed in [16].
We see that taken at face value the discussion in [21, 22] leads to the conclusion that “the recommended muonium hyperfine splitting” is the theoretical QED prediction for muonium HFS. There are too many reasons why this is not so. The estimate of the theoretical uncertainty $\delta_{\scriptscriptstyle th}(Mu)=-4(83)~Hz$ in Eq. (7) falls well inside the 85 Hz estimate of the yet uncalculated terms in this very text [21, 22] (or $70$ Hz as estimated in [16]). It is also hard to interpret how this $\delta_{\scriptscriptstyle th}(Mu)$ is compatible the $51$ Hz uncertainty on the right hand side in Eq. (7). In its turn this uncertainty on the right hand side in Eq. (7) is about 5 times lower than the uncertainly of the theoretical prediction in [20], and 10 times lower that the theoretical uncertainty in [16].

### v2

# Muonium Hyperfine Splitting Uncertainty Revisited

@@ Abstract
Uncertainty of the theoretical prediction for the hyperfine splitting in the ground state of muonium is considered. It is compared with the respective discussion in the two most recent CODATA adjustments of the fundamental physical constants.
High precision measurements of hyperfine splitting (HFS) in muonium for a long time were considered as a test of the high precision quantum electrodynamics (QED) and a source for precise values of the fine structure constant $\alpha$ and the muon-electron mass ratio $m_{\mu}/m_{e}$ [2, 3, 1]. While the role of muonium HFS in determining the fine structure constant was made obsolete by the highly precise $\alpha$ obtained from the measurements of the electron anomalous magnetic moment $a_{e}$ [4] and the recoil frequency of the 133Cs [5] and 87Rb [6] atoms, it remains the best source for the precise value of the muon-electron mass ratio. The hyperfine splitting in muonium and muon-electron mass ratio were extracted from the Zeeman splitting measurements some time ago in two LAMPF (Los Alamos Meson Physics Facility) experiments [7, 8]
$\Delta\nu^{ex}_{\scriptscriptstyle HFS}(\mbox{Mu})=4~463~302~776~(51)~\mbox{Hz},\quad\delta=1.1\times 10^{-8},$ (1)
@@
$\Delta\nu_{\scriptscriptstyle weak}$ is the $Z$-boson exchange contribution [9, 10, 11, 12], $\Delta\nu_{\scriptscriptstyle hadr}$ is the hadron vacuum polarization contribution [13, 14, 15, 16] and $\Delta\nu_{th}$ is the estimate of the yet uncalculated terms, see, e.g., [2, 3]. Numerically [17] [^1] This is an updated value of HFS, which takes into account some minor changes in theoretical contributions and values of the fundamental constants, which happened after [17] was published, see [17, 16, 18, 19, 15, 20]. Only the last digit changed by one, what does not change any conclusions of [17].
$\Delta\nu^{th}_{\scriptscriptstyle HFS}(\mbox{Mu})=4~463~302~873~(511)~(70)~(2)~\mbox{Hz},$ (5)
@@
$\Delta\nu^{th}_{\scriptscriptstyle HFS}(\mbox{Mu})=4~463~302~873~(515)~\mbox{Hz},\quad\delta=1.2\times 10^{-7}.$ (6)
@@
$\Delta\nu_{\scriptscriptstyle Mu}(th)+\delta_{\scriptscriptstyle th}(\mbox{Mu})=4~463~302~776~(51)~{\rm Hz}~~~~[1.1\times 10^{-8}]$ (7)
This “recommended value for the muonium hyperfine splitting” is never defined in the text but its value and uncertainty precisely coincide with the value and uncertainty of the experimental HFS in Eq. (1). This is not a surprise, if we assume that the experimental HFS was treated in [25, 26] as a fundamental constant and was included in the least square adjustment. As we already mentioned, the last precise experiment on muonium HFS was done in 1999 [7, 8], and up to this moment there are no comparable in accuracy sources for the experimental value of the muonium HFS. This means that no adjustment can significantly change the values and uncertainties of muonium HFS. Therefore, the very idea of including the experimental HFS in the least square adjustment seems to be dubious.
But this is not how “the recommended value for the muonium hyperfine splitting” is treated in [25, 26]. By appearance, “the recommended value for the muonium hyperfine splitting” [25, 26] in Eq. (7) looks like a theoretical QED prediction, it is called $\Delta E_{\scriptscriptstyle\mbox{Mu}}(\mbox{th})+\delta_{\scriptscriptstyle th}(\mbox{Mu})$. Moreover, “the recommended value for the muonium hyperfine splitting” is explicitly interpreted in [25, 26] as a theoretical QED prediction, it is compared with the theoretical QED prediction in [17]. The authors of [25, 26] write: “Eides (2019) gave an alternative prediction for the uncertainty of the recommended muonium hyperfine splitting.” This statement demonstrates that “the recommended muonium hyperfine splitting” is considered in [25, 26] as a theoretical QED prediction for muonium HFS since this is what was discussed in [17].
We see that taken at face value the discussion in [25, 26] leads to the conclusion that “the recommended muonium hyperfine splitting” is the theoretical QED prediction for muonium HFS. There are too many reasons why this is not so. The estimate of the theoretical uncertainty $\delta_{\scriptscriptstyle th}(\mbox{Mu})=-4(83)~\mbox{Hz}$ in Eq. (7) falls well inside the 85 Hz estimate of the yet uncalculated terms in this very text [25, 26] (or $70$ Hz as estimated in [17] [^3] The estimate of the yet uncalculated terms is based on the theoretical experience and can legitimately differ from one theorist to another. Anyway this minor discrepancy is irrelevant for the discussion in this paper.). It is also hard to interpret how this $\delta_{\scriptscriptstyle th}(\mbox{Mu})$ is compatible the $51$ Hz uncertainty on the right hand side in Eq. (7). In its turn this uncertainty on the right hand side in Eq. (7) is about 5 times lower than the uncertainly of the theoretical prediction in [24], and 10 times lower that the theoretical uncertainty in [17].